OpenStudy (pounce0129):

simplify each expression √44 - √11

OpenStudy (anonymous):

\[\sqrt{44} = \sqrt{4 \times 11} = \sqrt{4} \times \sqrt{11} \implies 2 \sqrt{11}\] So: \[2 \sqrt{11} - \sqrt{11} = ??\]